I think this is using the pythagorean theorem and I get really confused...
A mason wants to be sure he has a right angle corner of a building's foundation. He marks a point 3 ft from the corner along one wall and another point 4ft from the corner along the other wall. If the corner is a right angle, what should the distance be between the two marked points??
I did the a^2 + b^2 = c^2 formula and plugged in 3^2 + b^2 = 4^2 and got an -7 as an answer. I know a negative # can't be it so that's where I need assistance. Thanks in advance.

Answer by stanbon(75887)

A mason wants to be sure he has a right angle corner of a building's foundation. He marks a point 3 ft from the corner along one wall and another point 4ft from the corner along the other wall. If the corner is a right angle, what should the distance be between the two marked points??
----
"c" is the hypotenuse.
---
a^2 + b^2 = c^2
3^2 + 4^2 = c^2
9 + 16 = c^2
c^2 = 25
c = 5 ft,
-------------------------
Cheers,
Stan H.
===============